# The Open Journal of Astrophysics

### v1

@@ 7.3 Shared $f_{\rm dc}$ for iPTF16geu and SN Zwicky
If we make the assumption that the fraction of dark compact objects is the same across different galaxies, we can combine the constraints from iPTF16geu and SN Zwicky. This could be the case if the dark compact objects were all created in the very early Universe, and are spread evenly throughout the Universe and across galaxies. We performed another inference run with a joint Bayesian model, where the free parameters are $f_{\rm dc}$, $\eta$ Zwicky, $\log M_{*}$ Zwicky, $\mu$ Zwicky, $\eta$ iPTF16geu, $\log M_{*}$ iPTF16geu, and $\mu$ iPTF16geu. The data constraints we use are the same as in the individual fits. Fig. A.2 shows the resulting corner plots for all free parameters, where the results for $\eta$ are consistent with the individual fits. We obtain a tighter upper bound on the fraction of dark compact objects: $f_{\rm dc}$ $<0.19$ at the $95\%$ confidence level.
Figure 10: Lens galaxy mass slope, $\eta$, versus stellar mass, $M_{*}$, (left panel) and redshift (right panel), for iPTF16geu, SN Zwicky, and several lensed galaxies from the SLACS (Auger et al., 2010) and SL2S (Sonnenfeld et al., 2013) surveys. The lens galaxies of iPTF16geu and SN Zwicky have relatively low stellar masses (see section 5.1) and flat mass profiles (see section 7) compared to the population of observed lensed galaxies.
@@ 8 Discussion and conclusion
We developed a Bayesian model to simultaneously model our parameters of interest, taking into account constraints from observed quantities such as the stellar mass, magnification, lens galaxy light model, and SN image flux ratios. Our results show that there is no preference for additional dark compact objects in the lens galaxy, and we place $95\%$ upper limits of $f_{\rm dc}$ $<0.25$ for iPTF16geu and $f_{\rm dc}$ $<0.47$ for SN Zwicky. Our results are in agreement with $f_{\rm dc}$ measurements from strongly lensed quasars, which also find no evidence for any dark compact objects (e.g. Pooley et al., 2009, 2012; Schechter & Wambsganss, 2004; Schechter et al., 2014; Mediavilla et al., 2009, 2017; Esteban-Gutiérrez et al., 2020; Mediavilla & Jiménez-Vicente, 2024).
@@
We also constrain the mass slope of the lens galaxies, which is made possible because they are standardizable candles for which we know their magnification. We measure $\eta=1.70\pm 0.07$ for iPTF16geu and $\eta=1.81\pm 0.10$ for SN Zwicky. Compared to other known lens galaxies from SLACS and SL2S, these slope measurements are on the lower side, but not outside of what has been observed before. For SN Zwicky, which does not have a lensed host galaxy, this measurement using the SNIa intrinsic brightness is the only way to place tight constraints on the mass slope. For iPTF16geu, we can cross-check this measurement with the one obtained by fitting a PEMD lens model with the slope as a free parameter, as is done in Mörtsell et al. (2020), which is in agreement.
@@
Future work on microlensing of lensed SNe can take into account time varying microlensing effects, as well as a spectrum of different microlens masses. Additionally, in our work we did not consider the effects of mililensing, due to dark matter subhalos. The recent analysis of SN Zwicky by Larison et al. (2024), which included HST reference observations, investigated mililensing, and found its contribution to the total magnification to be negligible. Additionally, they investigated the flux ratio discrepancies of SN Zwicky in the $F475W$, $F625W$, $F814W$, and $F160W$ filters, confirming the evidence for microlensing and finding hints for a chromatic nature of the effects.
@@
This work demonstrates the power of strongly lensed SNe to obtain insights into compact objects and potential primordial black hole populations. With the advance of the next generation of telescopes such as the Vera C. Rubin Observatory (Ivezić et al., 2019) and the Nancy Grace Roman Space Telescope (e.g., Pierel et al., 2021), we will discover dozens more galaxy-scale lensed SNe, which will allow us to put tighter constraints on the population of dark compact objects. Moreover, these compact lensed SN systems offer one of the few avenues to probe the inner mass slope of lower-mass lens galaxies.
@@ Acknowledgments
We are grateful to Simon Birrer and Anowar Shajib for helpful suggestions regarding the lens modelling of iPTF16geu. We would like to thank Jorge Martin Camalich and Evencio Mediavilla for interesting and useful discussions, and Evencio Mediavilla for confirming our results with an independent re-analysis of the SN Zwicky data. This work has been enabled by support from the research project grant ‘Understanding the Dynamic Universe’ funded by the Knut and Alice Wallenberg Foundation under Dnr KAW 2018.0067. EM acknowledges support from the Swedish Research Council under Dnr VR 2020-03384. ST has been supported by funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programmes (grant agreement no. 101018897 CosmicExplorer). SD acknowledges a Kavli Fellowship and a JRF at Lucy Cavendish College. This project has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodovska-Curie grant agreement No 101105725. AG acknowledges support from the Swedish Research Council (Dnr 2020-03444) and the Swedish National Space Agency (Dnr 2023-00226). EEH is supported by a Gates Cambridge Scholarship (#OPP1144).
Software used: Astropy (Astropy Collaboration et al., 2013, 2018, 2022), Jupyter (Kluyver et al., 2016), Matplotlib (Hunter, 2007; Caswell et al., 2020), NumPy (Harris et al., 2020), Pandas (Wes McKinney, 2010; Team, 2023), Pickle (Van Rossum, 2020), SciPy (Virtanen et al., 2020), Seaborn (Waskom et al., 2020), Emcee (Foreman-Mackey et al., 2013; Goodman & Weare, 2010), ChainConsumer (Hinton, 2016), Lenstronomy (Birrer & Amara, 2018; Birrer et al., 2021), Prospector (Johnson et al., 2021, 2024), Dynesty (Speagle, 2020; Koposov et al., 2024), Reproject (Robitaille et al., 2020), SAOImageDS9 (Smithsonian Astrophysical Observatory, 2000).
@@
The simulation code for the figures and data analysis will be made publicly available on Github upon publication.
@@ Appendix A Appendix: corner plots
Fig. A.2 shows the resulting parameter posteriors when SN Zwicky and iPTF16geu are combined in one single fit, where they share a common $f_{\rm dc}$ parameter. By combining the two, a tighter upper limit of $f_{\rm dc}$ $<0.19\ (95\%)$ is obtained. This assumes that $f_{\rm dc}$ is the same at all image positions in both galaxies.

### v2

@@ 7.3 Shared $f_{\rm dc}$ for iPTF16geu and SN Zwicky
If we make the assumption that the fraction of dark compact objects is the same across different galaxies, we can combine the constraints from iPTF16geu and SN Zwicky. This could be the case if the dark compact objects were all created in the very early Universe, and are spread evenly throughout the Universe and across galaxies. We performed another inference run with a joint Bayesian model, where the free parameters are $f_{\rm dc}$, $\eta$ Zwicky, $\log M_{*}$ Zwicky, $\mu$ Zwicky, $\eta$ iPTF16geu, $\log M_{*}$ iPTF16geu, and $\mu$ iPTF16geu. The data constraints we use are the same as in the individual fits. Fig. A.2 shows the resulting corner plots for all free parameters, where the results for $\eta$ are consistent with the individual fits. We obtain a tighter upper bound on the fraction of dark compact objects: $f_{\rm dc}$ <0.19<0.19 at the $95\%$ confidence level.
Figure 10: Lens galaxy mass slope, $\eta$, versus stellar mass, $M_{*}$, (left panel) and redshift (right panel), for iPTF16geu, SN Zwicky, and several lensed galaxies from the SLACS (Auger2010) and SL2S (Sonnenfeld2013) surveys. The lens galaxies of iPTF16geu and SN Zwicky have relatively low stellar masses (see section 5.1) and flat mass profiles (see section 7) compared to the population of observed lensed galaxies.
@@ 8 Discussion and conclusion
We developed a Bayesian model to simultaneously model our parameters of interest, taking into account constraints from observed quantities such as the stellar mass, magnification, lens galaxy light model, and SN image flux ratios. Our results show that there is no preference for additional dark compact objects in the lens galaxy, and we place $95\%$ upper limits of $f_{\rm dc}$ <0.25<0.25 for iPTF16geu and $f_{\rm dc}$ <0.47<0.47 for SN Zwicky. Our results are in agreement with $f_{\rm dc}$ measurements from strongly lensed quasars, which also find no evidence for any dark compact objects (e.g. Pooley2009; Pooley2012; Schechter2004; Schechter2014; Mediavilla2009; Mediavilla2017; Esteban2020; Medavilla2024). With a larger sample of lensed SNe Ia, we will eventually be able to test whether the initial mass function (assumed to be Chabrier in this work) is correct, as we become more sensitive to $f_{\rm dc}$, or equivalently the stellar mass-to-light ratio (2025MNRAS.539..393W).
@@
We also constrain the mass slope of the lens galaxies, which is made possible because they are standardizable candles for which we know their magnification. We measure $\eta=1.70\pm 0.07$ for iPTF16geu and $\eta=1.81\pm 0.10$ for SN Zwicky. Compared to other known lens galaxies from SLACS and SL2S, these slope measurements are on the lower side, but not outside of what has been observed before. For SN Zwicky, which does not have a lensed host galaxy, this measurement using the SNIa intrinsic brightness is the only way to place tight constraints on the mass slope. For iPTF16geu, we can cross-check this measurement with the one obtained by fitting a PEMD lens model with the slope as a free parameter, as is done in Mortsell2020, which is in agreement.
We also investigated an alternative explanation for the observed flux ratio anomalies in iPTF16geu and SN Zwicky: angular complexity in the lens model. We tested lens models with fourth-order multipole perturbations and found that they cannot account for the observed flux ratios, leaving microlensing as the most likely cause.
@@
Future work on microlensing of lensed SNe can take into account time varying microlensing effects, as well as a spectrum of different microlens masses. Additionally, in our work we did not consider the effects of mililensing, due to dark matter subhalos. The recent analysis of SN Zwicky by Larison2024, which included HST reference observations, investigated mililensing, and found its contribution to the total magnification to be negligible. Additionally, they investigated the flux ratio discrepancies of SN Zwicky in the $F475W$, $F625W$, $F814W$, and $F160W$ filters, confirming the evidence for microlensing and finding hints for a chromatic nature of the effects.
@@
This work demonstrates the power of strongly lensed SNe to obtain insights into compact objects and potential primordial black hole populations. With the advance of the next generation of telescopes such as the Vera C. Rubin Observatory (ivezic2019lsst) and the Nancy Grace Roman Space Telescope (e.g., pierel2020projected), we will discover dozens more galaxy-scale lensed SNe, which will allow us to put tighter constraints on the population of dark compact objects. Moreover, these compact lensed SN systems offer one of the few avenues to probe the inner mass slope of lower-mass lens galaxies.
@@ Acknowledgments
We are grateful to Simon Birrer and Anowar Shajib for helpful suggestions regarding the lens modelling of iPTF16geu. We would like to thank Jorge Martin Camalich and Evencio Mediavilla for interesting and useful discussions, and Evencio Mediavilla for confirming our results with an independent re-analysis of the SN Zwicky data. We thank the referee for suggesting the inclusion of the multipole analysis in the paper.
This work has been enabled by support from the research project grant ‘Understanding the Dynamic Universe’ funded by the Knut and Alice Wallenberg Foundation under Dnr KAW 2018.0067. EM acknowledges support from the Swedish Research Council under Dnr VR 2020-03384. ST has been supported by funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programmes (grant agreement no. 101018897 CosmicExplorer). SD acknowledges funding by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding Guarantee EP/Z000475/1, a Kavli Fellowship and a JRF at Lucy Cavendish College. This project has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodovska-Curie grant agreement No 101105725. AG acknowledges support from the Swedish Research Council (Dnr 2020-03444) and the Swedish National Space Agency (Dnr 2023-00226). EEH is supported by a Gates Cambridge Scholarship (#OPP1144). This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (LensEra: grant agreement No. 945536). For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising.
Software used: AstroPy (Astropy2013; Astropy2018; Astropy2022), Jupyter (JupyterNotebook), Matplotlib (Matplotlib2007; Matplotlib2020), NumPy (Numpy202), Pandas (Pandas2010; Pandas2023), Pickle (Pickle2020), SciPy (Scipy2020), Seaborn (Seaborn2020), Emcee (ForemanMackey2013; Goodman2010), ChainConsumer (Hinton2016), Lenstronomy (birrer2018lenstronomy; Birrer2021LenstronomyII), Prospector (Johnson2021), FSPS (Conroy2009; Conroy2010; Conroy2010ApJ), Python-FSPS (Johnson2024), SEDPy (SEDPy2021), Dynesty (Speagle2020; Koposov2024), Reproject (Reproject2020), SAOImageDS9 (DS9; Joye2003).
@@
Code related to this article is available at https://github.com/Nikki1510/Microlensing
This paper was built using the Open Journal of Astrophysics template. The OJA is a journal which provides fast and easy peer review for new papers in the astro-ph section of the arXiv, making the reviewing process simpler for authors and referees alike. Learn more at http://astro.theoj.org.
@@ Appendix A Appendix: corner plots
Fig. A.2 shows the resulting parameter posteriors when SN Zwicky and iPTF16geu are combined in one single fit, where they share a common $f_{\rm dc}$ parameter. By combining the two, a tighter upper limit of $f_{\rm dc}$ $0.19\ (95\%)$ is obtained. This assumes that $f_{\rm dc}$ is the same at all image positions in both galaxies.